### v1

# Density matrix exponentiation and sample-based Hamiltonian simulation: Non-asymptotic analysis of sample complexity

Abstract
Density matrix exponentiation (DME) is a quantum algorithm that processes multiple copies of a program state $\sigma$ to realize the Hamiltonian evolution $e^{-i\sigma t}$. While serving as a prototypical sample-based quantum algorithm, DME is a powerful tool for various quantum information processing tasks, such as quantum principal component analysis and Hamiltonian simulation. In this work, we present a detailed sample complexity analysis of DME and sample-based Hamiltonian simulation. In particular, we prove that the sample complexity of DME is no larger than $4t^{2}/\varepsilon$, where $t$ is the desired evolution time and $\varepsilon$ is the desired imprecision level, as quantified by the normalized diamond distance. We also establish a fundamental lower bound on the sample complexity of sample-based Hamiltonian simulation, which matches our DME sample complexity bound up to a constant multiplicative factor, thereby proving that DME is optimal for sample-based Hamiltonian simulation. Finally, we point out that the DME sample complexity analysis in Appendix A of [Kimmel et al., npj Quantum Information 3, 13 (2017)] appears to be incomplete, highlighting the need for the results presented here, given the extensive use of DME over the past decade since its original proposal.
I Introduction
Density matrix exponentiation (DME) is a quantum algorithm that exponentiates a quantum state when multiple copies of it are available [1]. More specifically, given an unknown input quantum state $\rho$ and a sufficient number of copies of a “program” quantum state $\sigma$, DME approximately implements the unitary evolution $e^{-i\sigma t}$ on $\rho$ for a given evolution time $t$. DME is thus a particular method for achieving the task of sample-based Hamiltonian simulation [2], in which one realizes the unitary transformation $e^{-i\sigma t}$ approximately by consuming copies of a program state $\sigma$.
This exponentiation of quantum states offers a powerful tool for various quantum information processing tasks. As initially proposed in Ref. [1], it can be employed for quantum principal component analysis. Since the process of DME is independent of the particular states $\rho$ and $\sigma$, and the number of copies of the program state $\sigma$ for DME to achieve a desired precision level does not explicitly depend on its dimension, this gives an exponential advantage for quantum principal component analysis [1]. Also, DME offers a sample-based strategy for Hamiltonian simulation [2]. Specifically, if the Hamiltonian to be simulated can be encoded in a program state and multiple copies of it are available, one can simulate the dynamics induced by the corresponding Hamiltonian by using copies of the program state. Furthermore, DME can be employed in other tasks, such as block-encoding a density matrix when given access to its samples (as noted in [3] and investigated in [4, 5, 6]), quantum machine learning [1, 7], and revealing entanglement spectra [8]. Along with these applications, we have also witnessed experimental progress in implementing DME [9].
The essence of DME is that it becomes more accurate as the number of program states (i.e., sample number) increases. Here, a crucial question is to determine how many copies of the program state are required to realize DME up to the desired imprecision level, which is also referred to as the sample complexity of DME [2] (see also [10, 11] for the notion of sample complexity more generally). Refs. [1, 2] claimed that in the asymptotic limit, the sample complexity of DME is given by $O(t^{2}/\varepsilon)$, in order to achieve a desired evolution time $t\geq 0$ and imprecision level $\varepsilon\in[0,1]$. Ref. [2] also showed that this scaling behavior of the sample complexity is optimal, such that the sample complexity of an arbitrary protocol for sample-based Hamiltonian simulation cannot scale as $o(t^{2}/\varepsilon)$. However, as argued in Appendix A of our paper, the previous proof from [2, Appendix A], for the upper bound of DME sample complexity, is incomplete. That is, the proof given in [2, Appendix A] does not give a strict bound for an arbitrary evolution time.
In this paper, we rigorously establish the sample complexity of DME, in terms of the desired imprecision level $\varepsilon$ and evolution time $t$. More precisely, we show that the sample complexity of DME is no larger than $4t^{2}/\varepsilon$ (Theorem 2), where $t$ is the desired evolution time and $\varepsilon$ is the desired imprecision level, as quantified by the normalized diamond distance. We also investigate a fundamental lower bound on the sample complexity of an arbitrary protocol for sample-based Hamiltonian simulation by employing the notion of zero-error query complexity and appealing to prior results of [12] (see Theorem 5 and Lemma 7). This fundamental lower bound shows that the sample complexity of DME is optimal even in the non-asymptotic regime, up to a multiplicative constant factor.
Our paper is organized as follows. In Section II, we establish notation and review sample-based Hamiltonian simulation and DME. In Section III, we provide a formal definition of the sample complexity of sample-based Hamiltonian simulation. In Section IV, we establish an upper bound on the sample complexity of DME, and thus also on the sample complexity of sample-based Hamiltonian simulation. In Section V, we establish a lower bound on the sample complexity of sample-based Hamiltonian simulation, by means of the notion of zero-error query complexity of unitary channels. Finally, in Section VI, we provide some concluding remarks and directions for future research.
II Background
II.1 Notation
This section establishes some notation used throughout the rest of our paper. Let $\mathcal{H}_{S}$ denote the Hilbert space corresponding to a quantum system $S$. Let $\mathcal{D}(\mathcal{H}_{S})$ denote the set of quantum states (density operators) acting on the Hilbert space $\mathcal{H}_{S}$. We sometimes use the notation $\mathcal{H}^{d}$ to denote a Hilbert space in terms of its dimension $d$ and the notation $\mathcal{D}(\mathcal{H}^{d})$ to denote the set of $d$-dimensional quantum states.
[…]
Throughout this work, we use $\left\|\cdot\right\|_{1}$, which is referred to as the trace norm, and $\left\|\cdot\right\|_{\infty}$, which is referred to as the operator norm.
To quantify the distance (or closeness) between two quantum states $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{S})$, we use the normalized trace distance, which is equal to the normalized trace norm (i.e., $p=1$ in (1)) of their difference, such that
$\frac{1}{2}\left\|\rho-\sigma\right\|_{1}.$ (2)
This quantity is equal to the maximum difference in the probabilities that an arbitrary measurement operator can assign to these two states [13, Eq. (9.22)]. Note that the multiplicative factor $\frac{1}{2}$ above guarantees that $\frac{1}{2}\left\|\rho-\sigma\right\|_{1}\in[0,1]$ for any given quantum states $\rho$ and $\sigma$. Throughout this work, the normalization factor $\frac{1}{2}$ will sometimes be omitted when using the trace distance. We also use the fidelity of quantum states, which quantifies the closeness between two quantum states. Specifically, the fidelity $F(\rho,\sigma)$ between two quantum states $\rho,\sigma$ is defined as [14]
[…]
To quantify the distance between two quantum channels (completely positive and trace-preserving maps), we use the normalized diamond distance, which is defined for two quantum channels $\mathcal{N}$ and $\mathcal{M}$ as [15]
$\frac{1}{2}\left\|\mathcal{N}-\mathcal{M}\right\|_{\diamond}\coloneqq\frac{1}{2}\sup_{\rho\in\mathcal{D}\left(\mathcal{H}_{R}\otimes\mathcal{H}_{S}\right)}\left\|(\mathcal{I}_{R}\otimes\mathcal{N})(\rho)-(\mathcal{I}_{R}\otimes\mathcal{M})(\rho)\right\|_{1},$ (4)
where $R$ denotes a reference system with an arbitrarily large Hilbert space dimension and $\mathcal{I}_{R}$ denotes the identity channel acting on the reference Hilbert space $\mathcal{H}_{R}$. An important point is that, by definition, the dimension of $R$ can be arbitrarily large, but one can place a bound on the dimension of $R$ equal to the dimension of $S$ [16, Theorem 9.1.1]. Also, note that the normalized diamond distance between two quantum channels involves an optimization of the normalized trace distance between two arbitrary quantum states over $\mathcal{D}(\mathcal{H}_{R}\otimes\mathcal{H}_{S})$, and accordingly, it is guaranteed that $\frac{1}{2}\|\mathcal{N}-\mathcal{M}\|_{\diamond}\in[0,1]$, due to the normalization factor $\frac{1}{2}$.
[…]
$\operatorname{SWAP}_{S_{1}S_{2}}\coloneqq\sum_{i,j}|i\rangle\!\langle j|_{S_{1}}\otimes|j\rangle\!\langle i|_{S_{2}}.$ (5)
II.2 Review of sample-based Hamiltonian simulation and density matrix exponentiation
In this section, we review the task of sample-based Hamiltonian simulation and the DME algorithm for achieving this task [1, 2]. The aim of sample-based Hamiltonian simulation is to perform the following task: on input one copy of an unknown quantum state $\rho$ and $n$ copies of a program quantum state $\sigma$, implement the unitary operation $e^{-i\sigma t}$ on $\rho$ for evolution time $t\geq 0$ to within imprecision level $\varepsilon$ [1]. In short, the task is to perform a fixed quantum channel $\mathcal{P}^{(n)}$ (independent of the program state $\sigma$) such that the following inequality holds:
$\frac{1}{2}\left\|\mathcal{P}^{(n)}\!\left(\rho\otimes\sigma^{\otimes n}\right)-e^{-i\sigma t}\rho e^{i\sigma t}\right\|_{\diamond}\leq\varepsilon.$ (6)
Hereafter, we denote $\mathcal{U}_{\sigma,t}(\cdot)$ as the target unitary channel to be approximated by a sample-based Hamiltonian simulation algorithm, which is the ideal evolution according to the quantum state $\sigma$ for the evolution time $t\geq 0$. More formally,
$\mathcal{U}_{\sigma,t}(\rho)\coloneqq e^{-i\sigma t}\rho e^{i\sigma t}.$ (7)
We denote the number of copies of the input program state $\sigma$ by $n$, where $n\in\mathbb{N}$ and satisfies $n>t$.
One can divide the evolution time $t$ by the number of copies $n$, and we denote the resulting fraction as a unit time step $\Delta$; i.e., $\Delta\coloneqq\frac{t}{n}$. By means of the Hadamard lemma (see, e.g., [2, Lemma 19]), the ideal evolution of the state $\rho$ by the Hamiltonian $\sigma$ for a time step $\Delta$ can be expressed as a series in $\Delta$:
$\mathcal{U}_{\sigma,\Delta}(\rho)$ $=e^{-i\sigma\Delta}\rho e^{i\sigma\Delta}$ (8)
$=\rho-i\Delta[\sigma,\rho]-\frac{1}{2!}\Delta^{2}[\sigma,[\sigma,\rho]]+\cdots$ (9)
$=\sum_{j=0}^{\infty}\frac{(-i\Delta)^{j}}{j!}[(\sigma)^{j},\rho],$ (10)
where the nested commutator is defined as
$[(X)^{k},Y]\coloneqq\underbrace{[X,\dotsb[X,[X}_{k\text{ times }},Y]]\dotsb],\quad\quad[(X)^{0},Y]\coloneqq Y.$ (11)
Clearly, $n$ repetitions of the ideal evolution $\mathcal{U}_{\sigma,\Delta}$ lead to the desired ideal evolution $\mathcal{U}_{\sigma,t}$:
$\mathcal{U}_{\sigma,t}(\rho)=\mathcal{U}_{\sigma,\Delta}^{n}(\rho).$ (12)
The DME algorithm approximates the ideal time evolution $\mathcal{U}_{\sigma,\Delta}(\rho)$ for each time step $\Delta$ in (10) by utilizing a single copy of the program state $\sigma$. More specifically, given a quantum state $\rho$ in the system $S_{1}$ and a single copy of the program state $\sigma$ in the ancillary system $S_{2}$, DME applies the swap Hamiltonian between systems $S_{1}$ and $S_{2}$ for the unit time step $\Delta$, and then discards the ancillary system $S_{2}$. In short, for each step, DME realizes the following quantum channel:
$\widetilde{\mathcal{U}}_{\sigma,\Delta}(\rho_{S_{1}})\coloneqq\operatorname{Tr}_{S_{2}}\!\left[e^{-i\Delta\operatorname{SWAP}_{S_{1}S_{2}}}(\rho_{S_{1}}\otimes\sigma_{S_{2}})e^{i\Delta\operatorname{SWAP}_{S_{1}S_{2}}}\right],$ (13)
where $\operatorname{SWAP}_{S_{1}S_{2}}$ is defined in (5). For a sufficiently small time step $\Delta$, the quantum channel generated by DME in (13) is close to the ideal evolution in (10). More specifically, as argued in Ref. [1], the following holds:
$\operatorname{Tr}_{S_{2}}\!\left[e^{-i\Delta\operatorname{SWAP}_{S_{1}S_{2}}}(\rho_{S_{1}}\otimes\sigma_{S_{2}})e^{i\Delta\operatorname{SWAP}_{S_{1}S_{2}}}\right]=\rho-i\Delta[\sigma,\rho]+O(\Delta^{2}),$ (14)
where we neglected the system label $S_{1}$ on the right-hand side of (14) for more clarity. Since the right-hand side of (14) is equal to the right-hand side of (10) up to the first order in $\Delta$, the error of DME for each time step $\Delta$ is asymptotically bounded by the second order in $\Delta$, i.e.,
$\mathcal{U}_{\sigma,\Delta}(\rho)-\widetilde{\mathcal{U}}_{\sigma,\Delta}(\rho)=O(\Delta^{2}).$ (15)
Hence, for each step over $n$ steps, DME approximates the ideal evolution using a single copy of $\sigma$, up to the imprecision level $O(\Delta^{2})$. By repeating this process $n$ times (i.e., implementing $\widetilde{\mathcal{U}}_{\sigma,\Delta}^{n}(\rho)$) and thus consuming $n$ program states, DME finally approximates the ideal evolution $\mathcal{U}_{\sigma,t}(\rho)$.
It was argued in [1], by appealing to the Trotter–Suzuki theory of Hamiltonian simulation, that the total error scales as $O(n\Delta^{2})=O(t^{2}/n)$, so that $n=O(t^{2}/\varepsilon)$ copies of the program state are needed to have a total simulation error no larger than $\varepsilon$. The same conclusion was reached in [2] by means of a different argument. In fact, Ref. [2] claimed that for any input program state $\sigma$ the diamond distance between $\mathcal{U}_{\sigma,t}$ and $\widetilde{\mathcal{U}}_{\sigma,\Delta}^{n}$ is asymptotically bounded from above by $O(t^{2}/n)$. This indicates that $O(t^{2}/\varepsilon)$ samples are required to achieve the imprecision level $\varepsilon$. However, as argued in Appendix A, the proof of this previous sample complexity bound is incomplete, in the sense that it does not give a strict imprecision bound for an arbitrary evolution time $t$.
III Sample complexity of sample-based Hamiltonian simulation
The main goal of our paper is to examine the sample complexity of DME and sample-based Hamiltonian simulation. More formally, we investigate the precise number of program states, $n$, required to obtain a desired imprecision level $\varepsilon$ in the diamond distance for an evolution time $t$. As such, one of main goals for DME is to find the minimum value of $n$ such that the following error bound holds for arbitrarily given $t\geq 0$, $\varepsilon\in[0,1]$, and $\sigma$:
$\frac{1}{2}\left\|\mathcal{U}_{\sigma,t}-\widetilde{\mathcal{U}}_{\sigma,\Delta}^{n}\right\|_{\diamond}\leq\varepsilon.$ (16)
Let us now formally define the sample complexity, more generally, for sample-based Hamiltonian simulation, which includes the quantum channel induced by the DME algorithm. Here, the sample complexity is the minimum number of program states needed to approximate the unitary operation up to a desired imprecision level $\varepsilon$, where now the minimization is over all possible quantum channels.
Definition 1 (Sample complexity of sample-based Hamiltonian simulation)
The sample complexity of sample-based Hamiltonian simulation is denoted by $n_{d}^{\ast}(t,\varepsilon)$ and is defined as the minimum number of program states, the latter denoted by $\kappa$, required to realize a channel that is $\varepsilon$-distinguishable in normalized diamond distance from the ideal unitary evolution $\mathcal{U}_{\kappa,t}$, as defined in (7), for an arbitrary quantum state $\kappa$ of dimension $d$. Formally, the sample complexity $n_{d}^{\ast}(t,\varepsilon)$ is defined as
$n_{d}^{\ast}(t,\varepsilon)$ $\coloneqq\inf_{\mathcal{P}^{(n)}\in\rm{CPTP}}\left\{n\in\mathbb{N}:\frac{1}{2}\left\|\mathcal{P}^{(n)}\circ\mathcal{A}_{\kappa^{\otimes n}}-\mathcal{U}_{\kappa,t}\right\|_{\diamond}\leq\varepsilon\ \ \forall\kappa\in\mathcal{D}(\mathcal{H}^{d})\right\}$ (17)
$=\inf_{\mathcal{P}^{(n)}\in\rm{CPTP}}\left\{n\in\mathbb{N}:\sup_{\kappa\in\mathcal{D}(\mathcal{H}^{d})}\frac{1}{2}\left\|\mathcal{P}^{(n)}\circ\mathcal{A}_{\kappa^{\otimes n}}-\mathcal{U}_{\kappa,t}\right\|_{\diamond}\leq\varepsilon\right\},$ (18)
where $\mathcal{D}(\mathcal{H}^{d})$ denotes the set of quantum states of dimension $d$ and the appending channel $\mathcal{A}_{\kappa^{\otimes n}}$ for the quantum state $\kappa$ on an arbitrary input state $\zeta$ is defined as
$\mathcal{A}_{\kappa^{\otimes n}}(\zeta)\coloneqq\zeta\otimes\kappa^{\otimes n}.$ (19)
It is worth emphasizing that the quantum channel $\mathcal{P}^{(n)}$ in the right-hand side of (17) includes the DME operation. Therefore, an upper bound on the sample complexity of DME, with a program state being of dimension $d$, certainly provides an upper bound on $n_{d}^{\ast}$.
IV Upper bound on the sample complexity of density matrix exponentiation
In this section, we examine the sample complexity of DME (i.e., the number of program states needed to achieve a desired imprecision level $\varepsilon$ and evolution time $t$). This can be done by identifying an upper bound on the imprecision induced by the whole process of DME, in terms of sample number $n$ and evolution time $t$. Our main result here is stated in the following theorem:
Theorem 2
Let $t\geq 0$, let $n\in\mathbb{N}$ be such that $n>t$, and let $\Delta\coloneqq\frac{t}{n}$. For every quantum state $\sigma$, the error of density matrix exponentiation satisfies the following bound:
$\frac{1}{2}\left\|\mathcal{U}_{\sigma,t}-\widetilde{\mathcal{U}}_{\sigma,\Delta}^{n}\right\|_{\diamond}\leq\frac{4t^{2}}{n},$ (20)
where $\mathcal{U}_{\sigma,t}$ is defined in (7) and $\widetilde{\mathcal{U}}_{\sigma,\Delta}$ in (13).
Corollary 3 (Upper bound on the sample complexity of DME)
Theorem 2 implies the following upper bound on the sample complexity of sample-based Hamiltonian simulation:
$n_{d}^{\ast}(t,\varepsilon)\leq\frac{4t^{2}}{\varepsilon},$ (21)
for every dimension $d$.
Hence, Corollary 3 resolves the incompleteness of the claim from Ref. [2] and validates the previously claimed sample complexity bound $O(t^{2}/\varepsilon)$ from Ref. [1].
Proof of Theorem 2. Applying (12), we find that
$\frac{1}{2}\left\|\mathcal{U}_{\sigma,t}-\widetilde{\mathcal{U}}_{\sigma,\Delta}^{n}\right\|_{\diamond}$ $=\frac{1}{2}\left\|\mathcal{U}_{\sigma,\Delta}^{n}-\widetilde{\mathcal{U}}_{\sigma,\Delta}^{n}\right\|_{\diamond}$ (22)
$\leq n\cdot\frac{1}{2}\left\|\mathcal{U}_{\sigma,\Delta}-\widetilde{\mathcal{U}}_{\sigma,\Delta}\right\|_{\diamond}$ (23)
$\leq n(4\Delta^{2}),$ (24)
where we have inductively applied the subadditivity of the diamond distance (see Lemma 10) to obtain the first inequality. The second inequality follows from Lemma 4, which provides a precise upper bound on the error induced by DME, for each time step $\Delta$. The right-hand side of the second inequality then implies (20) after substituting $\Delta=\frac{t}{n}$.
In the rest of the section, we provide a precise error bound induced when simulating a single step of DME for the unit time $\Delta$.
Lemma 4
Let $\sigma$ be an arbitary quantum state, and suppose that $\Delta\in[0,1)$. For the quantum channels $\mathcal{U}_{\sigma,\Delta}$ and $\widetilde{\mathcal{U}}_{\sigma,\Delta}$ defined in (8) and (13), respectively, the following inequality holds:
$\frac{1}{2}\left\|\mathcal{U}_{\sigma,\Delta}-\widetilde{\mathcal{U}}_{\sigma,\Delta}\right\|_{\diamond}\leq 4\Delta^{2}.$ (25)
Proof. Let $\rho\in\mathcal{D}(\mathcal{H}_{R}\otimes\mathcal{H}_{S_{1}})$ be an unknown bipartite quantum state over the joint system $RS_{1}$, where $R$ is an arbitrary reference system. Also, let $\sigma\in\mathcal{D}(\mathcal{H}_{S_{2}})$ be a program quantum state over the system $S_{2}$. Then, from (13), we have
$(\mathcal{I}_{R}\otimes\widetilde{\mathcal{U}}_{\sigma,\Delta})(\rho_{RS_{1}})$
$=\operatorname{Tr}_{S_{2}}\!\left[(\mathbb{I}_{R}\otimes e^{-i\Delta\operatorname{SWAP}_{S_{1}S_{2}}})(\rho_{RS_{1}}\otimes\sigma_{S_{2}})(\mathbb{I}_{R}\otimes e^{i\Delta\operatorname{SWAP}_{S_{1}S_{2}}})\right]$ (26)
$=\operatorname{Tr}_{S_{2}}\!\left[\begin{array}[c]{c}\left(\mathbb{I}_{R}\otimes(\cos\Delta\cdot\mathbb{I}_{S_{1}S_{2}}-i\sin\Delta\cdot\operatorname{SWAP}_{S_{1}S_{2}})\right)(\rho_{RS_{1}}\otimes\sigma_{S_{2}})\\ \times\left(\mathbb{I}_{R}\otimes(\cos\Delta\cdot\mathbb{I}_{S_{1}S_{2}}+i\sin\Delta\cdot\operatorname{SWAP}_{S_{1}S_{2}}\right)\end{array}\right]$ (29)
$=\left(\cos^{2}\Delta\right)\operatorname{Tr}_{S_{2}}\!\left[(\rho_{RS_{1}}\otimes\sigma_{S_{2}})\right]+\left(\sin^{2}\Delta\right)\operatorname{Tr}_{S_{2}}\!\left[\operatorname{SWAP}_{S_{1}S_{2}}(\rho_{RS_{1}}\otimes\sigma_{S_{2}})\operatorname{SWAP}_{S_{1}S_{2}}\right]$
$\qquad-i\left(\sin\Delta\cos\Delta\right)\operatorname{Tr}_{S_{2}}\!\left[\operatorname{SWAP}_{S_{1}S_{2}}(\rho_{RS_{1}}\otimes\sigma_{S_{2}})-(\rho_{RS_{1}}\otimes\sigma_{S_{2}})\operatorname{SWAP}_{S_{1}S_{2}}\right]$ (30)
$=\left(\cos^{2}\Delta\right)\rho_{RS_{1}}-i\left(\sin\Delta\cos\Delta\right)\left[\left(\mathbb{I}_{R}\otimes\sigma_{S_{1}}\right),\rho_{RS_{1}}\right]+\left(\sin^{2}\Delta\right)\left(\operatorname{Tr}_{S_{1}}\left[\rho_{RS_{1}}\right]\otimes\sigma_{S_{1}}\right)$ (31)
$=\left(1-\sin^{2}\Delta\right)\rho_{RS_{1}}-i\frac{\sin 2\Delta}{2}\left[\left(\mathbb{I}_{R}\otimes\sigma_{S_{1}}\right),\rho_{RS_{1}}\right]+\left(\sin^{2}\Delta\right)\left(\operatorname{Tr}_{S_{1}}\left[\rho_{RS_{1}}\right]\otimes\sigma_{S_{1}}\right)$ (32)
$=\rho_{RS_{1}}-i\frac{\sin 2\Delta}{2}\left[\left(\mathbb{I}_{R}\otimes\sigma_{S_{1}}\right),\rho_{RS_{1}}\right]+\left(\sin^{2}\Delta\right)\left(\operatorname{Tr}_{S_{1}}\left[\rho_{RS_{1}}\right]\otimes\sigma_{S_{1}}-\rho_{RS_{1}}\right).$ (33)
where we used the fact that $e^{-i\Delta\operatorname{SWAP}_{S_{1}S_{2}}}=\cos\Delta\cdot\mathbb{I}_{S_{1}S_{2}}-i\sin\Delta\cdot\operatorname{SWAP}_{S_{1}S_{2}}$, given that the $\operatorname{SWAP}$ operator is self-inverse (i.e., $\operatorname{SWAP}^{2}=\mathbb{I}$). Also, from (10) and using the Hadamard Lemma in Ref. [2, Lemma 19], we have
$\left(\mathcal{I}_{R}\otimes\mathcal{U}_{\sigma,\Delta}\right)(\rho_{RS_{1}})$
$=\left(\mathbb{I}_{R}\otimes e^{-i\sigma\Delta}\right)\rho_{RS_{1}}\left(\mathbb{I}_{R}\otimes e^{i\sigma\Delta}\right)$ (34)
$=\rho_{RS_{1}}-i\Delta\left[\left(\mathbb{I}_{R}\otimes\sigma_{S_{1}}\right),\rho_{RS_{1}}\right]+\sum_{k=2}^{\infty}\frac{\left(-i\right)^{k}\Delta^{k}}{k!}\left[(\mathbb{I}_{R}\otimes\sigma_{S_{1}})^{k},\rho_{RS_{1}}\right],$ (35)
where we previously defined the nested commutator $\left[(\mathbb{I}_{R}\otimes\sigma_{S_{1}})^{k},\rho_{RS_{1}}\right]$ in (11). Then, the trace distance between $\left(\mathcal{I}_{R}\otimes\mathcal{U}_{\sigma,\Delta}\right)(\rho_{RS_{1}})$ and $(\mathcal{I}_{R}\otimes\widetilde{\mathcal{U}}_{\sigma,\Delta})(\rho_{RS_{1}})$ for an arbitrary quantum state $\rho_{RS_{1}}$ can be bounded from above as
$\left\|\left(\mathcal{I}_{R}\otimes\mathcal{U}_{\sigma,\Delta}\right)(\rho_{RS_{1}})-(\mathcal{I}_{R}\otimes\widetilde{\mathcal{U}}_{\sigma,\Delta})(\rho_{RS_{1}})\right\|_{1}$
$=\left\|\begin{array}[c]{c}-i\left(\Delta-\frac{\sin 2\Delta}{2}\right)\left[\left(\mathbb{I}_{R}\otimes\sigma_{S_{1}}\right),\rho_{RS_{1}}\right]+\sum_{k=2}^{\infty}\frac{\left(-i\right)^{k}\Delta^{k}}{k!}\left[(\mathbb{I}_{R}\otimes\sigma_{S_{1}})^{k},\rho_{RS_{1}}\right]\\ -\left(\sin^{2}\Delta\right)\left(\operatorname{Tr}_{S_{1}}\left[\rho_{RS_{1}}\right]\otimes\sigma_{S_{1}}-\rho_{RS_{1}}\right)\end{array}\right\|_{1}$ (38)
$\leq\left(\Delta-\frac{\sin 2\Delta}{2}\right)\left\|\left[\left(\mathbb{I}_{R}\otimes\sigma_{S_{1}}\right),\rho_{RS_{1}}\right]\right\|_{1}+\sum_{k=2}^{\infty}\frac{\Delta^{k}}{k!}\left\|\left[(\mathbb{I}_{R}\otimes\sigma_{S_{1}})^{k},\rho_{RS_{1}}\right]\right\|_{1}$
$\qquad+\left(\sin^{2}\Delta\right)\left\|\operatorname{Tr}_{S_{1}}\left[\rho_{RS_{1}}\right]\otimes\sigma_{S_{1}}-\rho_{RS_{1}}\right\|_{1}$ (39)
$\leq 2\left\|\sigma\right\|_{\infty}\left(\Delta-\frac{\sin 2\Delta}{2}\right)+\sum_{k=2}^{\infty}\frac{\left(2\Delta\left\|\sigma\right\|_{\infty}\right)^{k}}{k!}+2\sin^{2}\Delta$ (40)
$=2\left\|\sigma\right\|_{\infty}\left(\Delta-\frac{\sin 2\Delta}{2}\right)+e^{2\Delta\left\|\sigma\right\|_{\infty}}-1-2\Delta\left\|\sigma\right\|_{\infty}+2\sin^{2}\Delta$ (41)
$\leq\frac{4}{3}\left\|\sigma\right\|_{\infty}\Delta^{3}+\frac{9}{2}\Delta^{2}\left\|\sigma\right\|_{\infty}^{2}+2\Delta^{2}$ (42)
$\leq 8\Delta^{2},$ (43)
where we used in (40) the fact that
$\left\|\left[\left(\mathbb{I}_{R}\otimes\sigma_{S_{1}}\right),\rho_{RS_{1}}\right]\right\|_{1}\leq 2\left\|\mathbb{I}_{R}\otimes\sigma_{S_{1}}\right\|_{\infty}\left\|\rho_{RS_{1}}\right\|_{1}=2\left\|\sigma\right\|_{\infty},$ (44)
and then iterated this $k$ times to get that
$\left\|\left[\left(\mathbb{I}_{R}\otimes\sigma_{S_{1}}\right)^{k},\rho_{RS_{1}}\right]\right\|_{1}$ $\leq 2\left\|\sigma\right\|_{\infty}\left\|\left[\left(\mathbb{I}_{R}\otimes\sigma_{S_{1}}\right)^{k-1},\rho_{RS_{1}}\right]\right\|_{1}$ (45)
[…]
$\leq\left(2\left\|\sigma\right\|_{\infty}\right)^{k}\left\|\rho_{RS_{1}}\right\|_{1}$ (47)
$=\left(2\left\|\sigma\right\|_{\infty}\right)^{k}.$ (48)
We also used the fact that
$\left\|\operatorname{Tr}_{S_{1}}\left[\rho_{RS_{1}}\right]\otimes\sigma_{S_{1}}-\rho_{RS_{1}}\right\|_{1}\leq 2,$ (49)
given that both $\operatorname{Tr}_{S_{1}}\!\left[\rho_{RS_{1}}\right]\otimes\sigma_{S_{1}}$ and $\rho_{RS_{1}}$ are quantum states. We also used the following inequalities to establish (42):
$\Delta-\frac{\sin 2\Delta}{2}$ $\leq\frac{2}{3}\Delta^{3},$ (50)
$e^{2\Delta\left\|\sigma\right\|_{\infty}}-1-2\Delta\left\|\sigma\right\|_{\infty}$ $\leq\frac{9}{2}\Delta^{2}\left\|\sigma\right\|_{\infty}^{2},$ (51)
$\sin^{2}\Delta$ $\leq\Delta^{2},$ (52)
given that $\Delta<1$ and $\left\|\sigma\right\|_{\infty}\leq 1$.
Because the inequality in (43) holds for an arbitrary input state $\rho_{RS_{1}}$, the normalized diamond distance between $\mathcal{U}_{\sigma,\Delta}$ and $\widetilde{\mathcal{U}}_{\sigma,\Delta}$ is bounded from above by $4\Delta^{2}$, and we conclude (25).
V Lower bound on the sample complexity of sample-based Hamiltonian simulation
In this section, we derive a lower bound on the sample complexity of sample-based Hamiltonian simulation, which is consistent with the optimality result in Ref. [2]. Using the definition of the sample complexity given in Definition 1, the main result of this section is as follows:
Theorem 5
For all $t\geq 0$, $\varepsilon\in(0,a)$ where $a\coloneqq\min\left\{\frac{t}{8\pi e},\frac{79}{1000}\right\}$, and $d\in\{2,3,\ldots\}$, the following lower bound holds for the sample complexity of sample-based Hamiltonian simulation:
$n_{d}^{\ast}(t,\varepsilon)\geq\frac{4}{1000}\left(\frac{t^{2}}{\varepsilon}\right).$ (53)
The lower bound in Theorem 5 has the same scaling behavior as our sample complexity upper bound mentioned in Corollary 3 in terms of $t^{2}/\varepsilon$, although they have different multiplicative factors (4 versus $\frac{4}{1000}$). Hence, the lower bound in Theorem 5 matches the upper bound up to a multiplicative factor, thus establishing the sample complexity optimality of DME.
To show the lower bound on sample complexity, our strategy is to employ the notion of zero-error query complexity, which we define as the minimum number of copies required to perfectly distinguish the ideal unitary operation $e^{i\kappa t}$ for different quantum states $\kappa$. Before providing the proof of Theorem 5, we first introduce the zero-error query complexity and investigate a lower bound on the sample complexity of sample-based Hamiltonian simulation, in terms of this quantity. We then recover the lower bound on sample complexity in terms of $t$ and $\varepsilon$ as in (53).
V.1 Lower bound on the sample complexity of sample-based Hamiltonian simulation in terms of zero-error query complexity
In the following, we develop a sample complexity lower bound on sample-based Hamiltonian simulation, by employing the notion of zero-error query complexity. We first define the zero-error query complexity as the number of queries needed to distinguish two ideal unitary channels perfectly. Specifically, let $\mathcal{U}_{\kappa,t}(\cdot)=e^{-i\kappa t}(\cdot)e^{i\kappa t}$ be the ideal unitary evolution according to a quantum state $\kappa$, as we previously defined in (7). Then, the zero-error query complexity $m^{\ast}$ can be defined as follows:
Definition 6 (Zero-error query complexity [12])
Let $\rho,\sigma\in\mathcal{D}(\mathcal{H}^{d})$ be arbitrary $d$-dimensional quantum states such that $\rho\neq\sigma$. Then, we define the integer $m^{\ast}\equiv m^{\ast}(\rho,\sigma,t)\in\mathbb{N}$ as the minimum number of queries needed such that the two unitary channels $\mathcal{U}_{\rho,t}$ and $\mathcal{U}_{\sigma,t}$ for $t>0$ are perfectly distinguishable from each other. That is,
$m^{\ast}(\rho,\sigma,t)\coloneqq\inf\left\{m\in\mathbb{N}:\frac{1}{2}\left\|\mathcal{U}_{\rho,t}^{\otimes m}-\mathcal{U}_{\sigma,t}^{\otimes m}\right\|_{\diamond}=1\right\}.$ (54)
It is worth emphasizing that, for arbitrary $\mathcal{U}_{\rho,t}$ and $\mathcal{U}_{\sigma,t}$ such that $t>0$ and $\rho\neq\sigma$, the value $m^{\ast}$ is finite, as shown in Ref. [12].
Hereafter, we derive a lower bound on the sample complexity of sample-based Hamiltonian simulation, using the zero-error query complexity $m^{\ast}$ that we have just defined.
Lemma 7
Let $\rho,\sigma\in\mathcal{D}(\mathcal{H}^{d})$ be arbitrary quantum states such that $\rho\neq\sigma$. Then we have the following lower bound on the sample complexity $n_{d}^{\ast}(t,\varepsilon)$:
$n_{d}^{\ast}(t,\varepsilon)\geq\frac{\ln\left[4\varepsilon\,m^{\ast}(\rho,\sigma,t)\right]}{m^{\ast}(\rho,\sigma,t)\ln F(\rho,\sigma)},$ (55)
where $F(\rho,\sigma)$ is the quantum fidelity between $\rho$ and $\sigma$, defined in (3).
Remark 8
Observe that the lower bound in (55) is non-negative (and thus non-trivial) only when $14\varepsilon\,m^{\ast}(\rho,\sigma,t)<$ 1, or equivalently when $\varepsilon<\frac{1}{4m^{\ast}(\rho,\sigma,t)}$.
Proof of Lemma 7. We first loosen the condition of $n_{d}^{\ast}(t,\varepsilon)$ in (17). Specifically, instead of considering all possible quantum states $\kappa\in\mathcal{D}(\mathcal{H}^{d})$ on the right-hand side of (17), we only consider two fixed quantum states $\rho,\sigma\in\mathcal{D}(\mathcal{H}^{d})$ such that $\rho\neq\sigma$. For these states, let us define $n^{\ast}(\rho,\sigma,t,\varepsilon)$ to be equal to the minimum number of samples of these states needed to realize a channel that is $\varepsilon$-distinguishable in the diamond distance from the ideal unitary evolution:
$n^{\ast}_{d}(\rho,\sigma,t,\varepsilon)\coloneqq\inf_{\mathcal{P}^{(n)}\in\text{CPTP}}\left\{\begin{array}[c]{c}n\in\mathbb{N}:\frac{1}{2}\left\|\mathcal{P}^{(n)}\circ\mathcal{A}_{\rho^{\otimes n}}-\mathcal{U}_{\rho,t}\right\|_{\diamond}\leq\varepsilon,\\ \qquad\quad\,\frac{1}{2}\left\|\mathcal{P}^{(n)}\circ\mathcal{A}_{\sigma^{\otimes n}}-\mathcal{U}_{\sigma,t}\right\|_{\diamond}\leq\varepsilon\end{array}\right\}.$ (56)
Then, for all quantum states $\rho,\sigma\in\mathcal{D}(\mathcal{H}^{d})$, we have that
$n_{d}^{\ast}(t,\varepsilon)\geq n_{d}^{\ast}(\rho,\sigma,t,\varepsilon),$ (57)
because the right-hand side of (56) is a relaxation of the right-hand side of (17). Accordingly, our revised goal here is to find a lower bound on $n_{d}^{\ast}(\rho,\sigma,t,\varepsilon)$ in terms of $m^{\ast}(\rho,\sigma,t)$, which will in turn serve as a lower bound on $n_{d}^{\ast}(t,\varepsilon)$ by the inequality in (57).
[…]